OpenStudy (anonymous):

How do I subtract 2pi from 7?

OpenStudy (anonymous):

I still can't subtract that

OpenStudy (anonymous):

then I end up w./ 7\[7\pi/\pi- 2\pi^(2)/\pi\]

OpenStudy (anonymous):

that's one way, I was looking for the angle 7 rad within the domain (-pi/2,pi/2)

OpenStudy (anonymous):

w/ decimals you end up w/ 0.72